Prove that if n is an integer and 3n + 2 is even, then n is even using

a) a proof by contraposition.

b) a proof by contradiction


1
Expert's answer
2021-07-13T08:53:05-0400

a) Assume the negation of conclusion:

Let "n" is an integer. Assume that "n" is odd. By definition of odd numbers, exists integer "k" such that "n=2k+1."

Hence


"3n+2=3(2k+1)+2=6k+5=2(3k+2)+1"

So by definition of odd numbers, "3n+2" is odd number. Thus we prove the negation of hypothesis.

Therefore we prove that if "n" is an integer and "3n+2" is even, then "n" is even using a proof by contraposition.


b)

Let "n" is an integer. Assume that "3n+2" is even and "n" is odd. Then by definition of odd numbers, there exists an integer "k" such that "n=2k+1." Hence


"3n+2=3(2k+1)+2=6k+5=2(3k+2)+1"

So by definition of odd numbers, "3n+2" is odd number, that contradicts our assumption that "3n+2" is even.

Hence, it is not the case that "3n+2" is even and "n" is odd.

Therefore we prove that if "n" is an integer and "3n+2" is even, then "n" is even using a proof by contradiction.